OpenStudy (anonymous):

determine the probability of e) μ = 37, σ = 4.35, x > 35: f) μ = 156, σ = 11.4, x ≥ 170: round z to 2 decimals. round answer to 4 decimal.

OpenStudy (amistre64):

how to find z?

OpenStudy (amistre64):

z is the standard score that makes all normal distributions stand upon equal footings

OpenStudy (amistre64):

z is a measure of how far away you are from a mean of 0 with a standard deviation of 1

OpenStudy (amistre64):

if the mean is 37 and we want to get it to 0, we have to subtract ...well, 37 from it and all its data points; we shift it to 0

OpenStudy (amistre64):

(x - mean) is a good start, but we need to adjust this for the standard deviation not; it needs to be 1. And to do that you simply divide the given sd by itself ... and all points associsated with it

OpenStudy (amistre64):

z = (x-mean)/sd

OpenStudy (amistre64):

to find the zscore for say : 35 then.. z = (35-37)/4.35 = 2/4.35 = .459

OpenStudy (amistre64):

now we need to either integrate the equation of the normal curve distribution ... aint gonna happen; or use tables that were created from them to look up this z scores probability value

OpenStudy (amistre64):

the trick is being able to read the ztables since there are a few different ones that measure from different points

OpenStudy (amistre64):

let me find one I can use from google tho ...this one should do fine

OpenStudy (anonymous):

I just figured this one out can you help me with the one i just posted?

OpenStudy (amistre64):

this one measures from all the way to the left so it gives us the probability value for the area LESS THAN our z

OpenStudy (amistre64):

lol ... i can try :)